OpenStudy (anonymous):

Ok now a tough one ! 2 people standing on a rotating disk in angular velocity omega. distance between them is 2d and the center of the disk is between them. one is sliding a box to the other person, the other person catches it after time T. Find the initial velocity vector of the box...

OpenStudy (mani_jha):

v(box w.r.t. disk) = v(box)-v(disk) 2d/T=v(box)-wd v(box)=2d/T + wd Is that correct?

OpenStudy (anonymous):

i agree with mani jha but this is not 1 dimensional motion so the angular velocity of the box d.w and the linear velocity of box (2d/T) need not act in the same direction right???so we may have to add them vectorially......

OpenStudy (mani_jha):

yeah, u r right, 2d/T is along the radial direction and w.d is along the tangential. So the correct answer would be (2d/T)i+(wd)j